Simplify and express each of the following in exponential form:
 (a^5 /a^3 ) × a^8

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

= (a^5- 3) × a^8 ... [∵a^m ÷a^n = a^m-n ]
= a^2× a^8
= a^2+8 ... [∵a^m × a^n = a^m+n ]
= a^10

​Thanks